Log(5x+2)-log(x-1)=0.7782. Find x
THANK YOU

Answer by fractalier(6550)

Okay, from
log(5x+2)-log(x-1)=0.7782
combine logs first and get
log((5x+2)/(x-1)) = .7782
Now exponentiate 10-to-the and get
(5x+2)/(x-1) = 10^.7782
(5x+2)/(x-1) = 6
Now cross-multiply
5x + 2 = 6x - 6
and solve, giving us
x = 8